# ALMAGAL VIII. Cataloging Hierarchical Mass Structure from Cores to Clumps across the Galactic Disk

### v1

@@ 3.3 Calculation of Continuum Structure Properties
where $d$ is the distance to the ALMAGAL source from Molinari et al. (2025), $F_{\nu}$ is the integrated flux of the structure, $R_{gd}$ is the gas-to-dust ratio, $\kappa_{\nu}$ is the dust opacity per unit mass, and $B_{\nu}(T_{d})$ is the Planck function at the local dust temperature $T_{d}$. We take the gas-to-dust ratio to be 100 and we use a dust opacity per unit mass of $\kappa_{\text{1.3mm}}=0.9$~cm2g-1 based on the Ossenkopf & Henning (1994) model for moderately coagulated dust grains with thin ice mantles at gas densities $\geq 10^{6}$cm-3. We also calculate a background subtracted mass using the background subtracted integrated flux of the structure, $F_{\nu,bgs}$.
@@
The surrounding surface density $\Sigma_{\text{surr}}$ is calculated by subtracting the total mass of cores, $M_{\text{cores}}$, contained within a dendrogram structure from the mass of the dendrogram structure, $M_{\text{struc}}$
@@
Figure 4: Violin plots showing the distribution of heliocentric distances (Top) and Galactocentric radii (Middle) and the clump mass (Bottom) for the different morphological categories of clumps in the ALMAGAL sample (Empty, Isolated, Simple, Rich and All). The center vertical line indicates the median for each distribution.
4 Results
We present the distributions of various clump properties as they relate to the hierarchical morphology observed in our ALMA data. We look at 4 subcategories in our sample covering different levels of fragmentation: Empty, Isolated, Simple, and Rich. The Isolated, Simple, and Rich targets are defined based on the method outlined in Section 3.2.
@@ 4 Results
4.1 Evaluating Distance Bias
@@
4.2 Comparing Hierarchical Morphology and Clump Properties
In Figure 5 we present violin plots showing the distributions of clump physical properties for the full sample as well as for the different morphological sub-samples we define in Section 3.2. We find that targets with more complex hierarchical structure tend to have distributions containing higher surface densities, warmer dust temperatures, and larger L/M ratios at the clump-scale. The Rich sample has a median mass surface density of 1.28 g cm-2, the highest out of the different morphological sub-categories. The median surface density continues to decrease as we move from the Rich to the Empty sub-samples, where Empty targets have a median surface density of 0.35 g cm-2.
The Rich sub-sample contains the clumps with the highest dust temperatures (up to 39.5 K), and have a median $T_{\text{dust}}$ $\sim 24$ K. This is a $\geq 10$ K offset from the median $T_{\text{dust}}$ of the Empty and Isolated sub-samples, with values of 12.2-13.2 K, respectively.
Figure 5: Violin plots highlighting the population differences in clump surface density (Top Left), dust temperature (Top Right), L/M ratio (Bottom Left), and MMC mass (Bottom Right) across the Empty, Isolated, Simple, and Rich clump sub-samples as well as the full sample. The center vertical line indicates the median for each distribution. In general, the morphologically rich targets have a higher median surface density, dust temperature, MMC mass, and $L/M$, all of which are properties that may be related to the evolutionary stage of the clump.
@@ 4.2 Comparing Hierarchical Morphology and Clump Properties
4.3 K-S Tests
Figure 6: Heatmap matrix for Two Sample K-S Test p-values for each combination of the Rich (R), Simple (S), Isolated (I), and Empty (E) source populations. All p-values are shown in white text. A p-value =0.05=0.05 is shown with a solid black line on the color bar. The p-values <0.05<0.05 indicate combinations where we reject the null hypothesis that both samples are pulled from an identical distribution at a confidence level of 95%. The clump properties are listed on the $y$-axis, where $d_{\text{clump}}$ is the heliocentric distance, $R_{\text{GCR}}$ is the Galactocentric radius, $L_{\text{clump}}$ is the clump luminosity, $M_{\text{clump}}$ is the clump mass, $L/M$ is the clump luminosity-to-mass ratio, $\Sigma_{\text{clump}}$ is the clump surface density, $D_{\text{clump}}$ is the clump diameter, $n_{\text{frag}}$ is the fragment density of cores from the ALMAGAL core catalog (Coletta et al., 2025) in the clump, and $T_{\text{clump}}$ is the clump dust temperature.
We quantitatively determine if the clump property distributions for the different sub-categories (Rich, Simple, Isolated, Empty) are statistically distinct using a two sample Kolmogorov-Smirnov (K-S) test. The null hypothesis for this test is that the samples from the two data sets come from the same underlying distribution. The K-S test produces a probability values (p-value) that is compared to a chosen significance level (most commonly 0.05). We use the kstest function from the Python package scipy. A summary of the K-S test of each clump property for each sub-population category is given in Table 6.
Most sub-sample combinations have statistically similar distributions for heliocentric distance and Galactocentric radius. The exception to this is the Empty clump sub-sample, which has distance and Galactocentric radius distributions that are distinct from the other sub-samples, likely because of the lower dendrogram detection rate in ‘Far’ sample sources as discussed in Section 4.1. We also note that the Rich and Simple distributions have statistically dissimilar distributions for heliocentric distance at the 95% confidence level. This is likely due to rounding our maximum dendrogram level thresholds to the nearest integer for our morphological classifications, which causes a lack of precision in ensuring a perfectly even sampling of morphological groups between the near and far sample.
@@ 4.4 Structure Properties
In the right panel of Figure 7, we see how the mass and radii of our structures vary as a function of their morphological sub-category. All three groups of targets follow the same trend, with regions exhibiting more hierarchical structure containing the largest and most massive structures. For the Rich, Simple, and Isolated sub-categories, the median masses of structures are $\sim$[6.8, 4.1, 2.0] $M_{\odot}$, and the median effective radii are $\sim$[4290, 3940, 3150] au, respectively.
In Appendix D, we evaluate structure property differences between the near and far sample, and find that both samples extract structures of similar size and mass. We note that the near sample observations have more low-mass structures, but this is to be expected, as the far sample has a poorer mass sensitivity.
@@
5 Discussion
5.1 Hierarchical Structure and Clump Evolution
@@
In Section 3.2, we examined various clump evolutionary indicators, such as dust temperature, surface density, and MMC mass, and compared their distributions for our different morphological sub-samples (Empty, Isolated, Simple, and Rich). In general, we find that clumps exhibiting an increased level of hierarchical complexity have distributions shifted towards higher dust temperatures, surface densities and MMC masses. The KS-tests performed in Section 4.3 confirm that our morphological sub-samples have distinct distributions in most of the properties surveyed, with significant ($p<0.05$) evidence they do not draw from the same underlying distribution. The Rich sub-sample of ALMAGAL targets, which show the most structural hierarchy, have an L/M ratio distribution that heavily skews towards higher L/M, with a median value of 20.7 L/M. We consider clumps with L/M >10>10 L/M to be at a later stage of evolution, undergoing active star formation with the possible presence of Hii regions (Cesaroni et al., 2025).
@@
Additionally, as seen in Figure 5, the median MMC masses, dust temperatures and deconvolved surface densities for each clump increases as regions exhibit more morphological complexity. This may be due to the increasing fragmentation of the clump as it evolves (Coletta et al., 2025). Additionally, the amount of material being funneled down to the size scales that we can observe in the ALMAGAL observations should be higher for more evolved regions, and may contribute to the increased richness of structure that we see towards those targets.
5.2 Core Mass Accretion at Different Evolutionary Stages
@@
We investigate the correlation between the MMC mass in a clump and its surrounding surface density, $\Sigma_{\text{surr}}$ (see Section 3.3 for details) in the left panel of Figure 8. We see that the MMC mass is directly related to the surrounding surface density in its environment, with MMCs existing only in proportionally dense environments. Different evolutionary stages follow a similar linear trend, with clumps at later evolutionary stages (higher $L/M$) exhibiting more massive MMCs in higher density environments.
We extend this analysis of comparing the core mass with its surrounding surface density, $\Sigma_{\text{surr}}$, to all sources in the ALMAGAL core catalog for which there is a corresponding dendrogram structure, and present these results in the right panel of Figure 8.
@@
When we include the entire core population, we find that the core mass has a looser correlation with its surrounding surface density than when we consider just the MMCs. For all three evolutionary populations shown in Figure 8, the distribution of $M_{\text{core}}$ vs $\Sigma_{\text{surr}}$ includes a larger parameter space than the distribution of $M_{\text{MMC}}$ vs $\Sigma_{\text{surr}}$, with the inclusion of lower mass cores located in higher surrounding surface density environments. This spread in the distribution increases for clumps with a higher $L/M$ ratio.
@@
One interpretation is that cores competitively accrete mass from an intra-clump reservoir of gas, with some cores accreting mass more efficiently than others depending on their location within the clump (Bonnell et al., 2001; Bonnell & Bate, 2006). In this competitive accretion scenario, cores that are formed later, or in a location outside of the central region of the clump will have a slower mass accretion rate, even as more gas is funneled down from the larger scale reservoir. If we examine the right panel of Figure 8 in the context of core mass accretion throughout the evolution of a clump, we see that cores in clumps at early stages of their evolution have a tighter relationship between $M_{\text{core}}$ and $\Sigma_{\text{surr}}$. As this clump evolves, certain cores will continue to accrete material and maintain a $\Sigma_{\text{surr}}$ that is proportional to their mass. This is especially true for the MMCs, which have masses that are proportional to their surrounding surface densities, regardless of clump evolutionary stage. However in the same amount of time, other member cores will accrete less mass, while a steady stream of material from the intra-clump reservoir continues to increase their surrounding surface densities, which would reproduce the weaker $M_{\text{core}}$ vs. $\Sigma_{\text{surr}}$ correlation that we observe for clumps with higher $L/M$ ratios.
@@
5.3 Sources of Uncertainty
We acknowledge that there are multiple sources of uncertainty in our measurements, in particular our mass estimations as they are highly reliant on our dust temperature assumptions. However, as discussed in Appendix B, if we instead assume the dust temperature of the structure is equal to the dust temperature of its host clump, $T_{\text{clump}}$, we find that the relationship shown in Figure 8 still holds, but exhibits a systematic shift towards higher surrounding surface densities. This makes sense as $T_{\text{clump}}$ is often less than the assumed core-scale temperature from the model we use in Section 3.3, which would result in higher masses, and consequently, surface densities for all structures.
@@ 5.3 Sources of Uncertainty
We also explored how missing flux caused by short-spacings could impact our results. In Appendix C, we discuss single-dish combination using continuum observations from the ATLASGAL Survey (Schuller, F. et al., 2009), and quantitatively determine that for most structures, mass calculations differ by less than 10% when using the single-dish combined data instead. Additionally, since feathered data includes more extended emission, our dendrogram structures would have higher measured surface densities while the core masses would remain largely the same. In this scenario, the range of local surface densities calculated from dendrogram structures would increase, causing a larger spread along the $x$-axis of Figure 8, which is still consistent with the results we present in Section 5.2. Given this information, and the fact that only 856 out of 900 targets have overlap with the ATLASGAL survey, we ultimately decided against using single-dish combined images in our analysis.
@@ 6 Conclusion
    Using astrodendro, we were able to identify 5160 structures across 904 dense clumps targeted in the ALMAGAL survey. These structures have effective radii spanning 80042000800-42000 au and cover multiple orders of magnitude in mass, from 0.056700.05-670 M.
@@
In this paper, we described the process for generating the ALMAGAL hierachical continuum structure catalog and presented the distribution of the fundamental properties of these structures. We found that the level of fragmented sub-structure in clumps is related to their evolution. Additionally, we used the hierarchical continuum structure to infer the mass evolution of cores in clumps at different evolutionary stages, finding evidence of a clump-fed accretion scenario.
7 Acknowledgments

### v2

@@ III.3 Calculation of Continuum Structure Properties
where $d$ is the distance to the ALMAGAL source from Molinari et al. (2025), $F_{\nu}$ is the integrated flux of the structure, $R_{gd}$ is the gas-to-dust ratio, $\kappa_{\nu}$ is the dust opacity per unit mass, and $B_{\nu}(T_{d})$ is the Planck function at the local dust temperature $T_{d}$. We take the gas-to-dust ratio to be 100 and we use a dust opacity per unit mass of $\kappa_{\text{1.3mm}}=0.9$~cm2g-1 based on the Ossenkopf and Henning (1994) model for moderately coagulated dust grains with thin ice mantles at gas densities $\geq 10^{6}$cm-3 and a coagulation timescale of 105 yr. We also calculate a background subtracted mass using the background subtracted integrated flux of the structure, $F_{\nu,bgs}$.
@@
The surrounding surface density $\Sigma_{\text{surr}}$ is calculated by subtracting the total non-background subtracted mass of cores, $M_{\text{cores}}$, contained within a dendrogram structure from the mass of the dendrogram structure, $M_{\text{struc}}$
@@
Figure 4: Violin plots showing the distribution of heliocentric distances (Top) and Galactocentric radii (Bottom) for the different morphological categories of clumps in the ALMAGAL sample (Empty, Isolated, Simple, Rich and All). The center vertical line indicates the median for each distribution.
IV Results
We present the distributions of various clump properties as they relate to the hierarchical morphology observed in our ALMA data. We look at 4 subcategories in our sample covering different levels of fragmentation: Empty, Isolated, Simple, and Rich. The Isolated, Simple, and Rich targets are defined based on the method outlined in Section III.2.
@@ IV Results
IV.1 Evaluating Distance Bias
@@
IV.2 Comparing Hierarchical Morphology and Clump Properties
In Figure 5 we present violin plots showing the distributions of clump physical properties for the full sample as well as for the different morphological sub-samples we define in Section III.2. We find that targets with more complex hierarchical structure tend to have distributions containing higher surface densities, warmer dust temperatures, and larger L/M ratios at the clump-scale. The Rich sample has a median mass surface density of 1.28 g cm-2, the highest out of the different morphological sub-categories. The median surface density continues to decrease as we move from the Rich to the Empty sub-samples, where Empty targets have a median surface density of 0.35 g cm-2.
The Rich sub-sample contains the clumps with the highest dust temperatures (up to 39.5 K), with a median $T_{\text{dust}}$ $\sim 24$ K. This is a $\geq 10$ K offset from the median $T_{\text{dust}}$ of the Empty and Isolated sub-samples, with values of 12.2-13.2 K, respectively.
Figure 5: Violin plots highlighting the population differences in clump surface density (Top Left), dust temperature (Top Right), mass ($M_{\text{clump}}$) (Middle Left) L/M ratio (Middle Right), MMC mass (Bottom Left), and core number density ($n_{\text{frag}}$) across the Empty, Isolated, Simple, and Rich clump sub-samples as well as the full sample. The center vertical line indicates the median for each distribution. In general, the morphologically rich targets have a higher median surface density, dust temperature, L/M, MMC mass, and core number density, all of which are properties that may be related to the evolutionary stage of the clump. The mass is the only clump property that does not vary significantly between the different mophological subsamples.
@@ IV.2 Comparing Hierarchical Morphology and Clump Properties
IV.3 K-S Tests
Figure 6: Heatmap matrix for Two Sample K-S Test p-values for each combination of the Rich (R), Simple (S), Isolated (I), and Empty (E) source populations. All p-values are shown in white text. A p-value =0.05=0.05 is shown with a solid black line on the color bar. The p-values <0.05<0.05 indicate combinations where we reject the null hypothesis that both samples are pulled from an identical distribution at a confidence level of 95%. The clump properties are listed on the $y$-axis, where $d_{\text{clump}}$ is the heliocentric distance, $R_{\text{GCR}}$ is the Galactocentric radius, $L_{\text{clump}}$ is the clump luminosity, $M_{\text{clump}}$ is the clump mass, $L/M$ is the clump luminosity-to-mass ratio, $\Sigma_{\text{clump}}$ is the clump surface density, $D_{\text{clump}}$ is the clump diameter, $n_{\text{frag}}$ is the fragment density of cores from the ALMAGAL core catalog (Coletta et al., 2025) in the clump (the number of cores divided by the area of the observed field of view), and $T_{\text{clump}}$ is the clump dust temperature.
We quantitatively determine if the clump property distributions for the different sub-categories (Rich, Simple, Isolated, Empty) are statistically distinct using a two sample Kolmogorov-Smirnov (K-S) test. The null hypothesis for this test is that the samples from the two data sets come from the same underlying distribution. The K-S test produces a probability values (p-value) that is compared to a chosen significance level (most commonly 0.05). We use the kstest function from the Python package scipy. A summary of the K-S test of each clump property for each sub-population category is given in Figure 6.
Most sub-sample combinations have statistically similar distributions for heliocentric distance and Galactocentric radius. The exception to this is the Empty clump sub-sample, which has distance and Galactocentric radius distributions that are distinct from the other sub-samples, likely because of the lower dendrogram detection rate in ‘Far’ sample sources as discussed in Section IV.1. We also note that the Rich and Simple distributions have statistically dissimilar distributions for heliocentric distance at the 95% confidence level. This is likely due to rounding our maximum dendrogram level thresholds to the nearest integer for our morphological classifications, which causes a lack of precision in ensuring a perfectly even sampling of morphological groups between the near and far sample.
@@ IV.4 Structure Properties
In the right panel of Figure 7, we show how the mass and radii of our structures vary as a function of their morphological sub-category. All three groups of targets follow the same trend, with regions exhibiting more hierarchical structure containing the largest and most massive structures. For the Rich, Simple, and Isolated sub-categories, the median masses of structures are $\sim$[6.8, 4.1, 2.0] $M_{\odot}$, and the median effective radii are $\sim$[4290, 3940, 3150] au, respectively.
In Appendix E, we evaluate structure property differences between the near and far sample, and find that both samples extract structures of similar size and mass. We note that the near sample observations have more low-mass structures, but this is to be expected, as the far sample has a poorer mass sensitivity.
@@
V Discussion
V.1 Hierarchical Structure and Clump Evolution
@@
In Section III.2, we examined various clump evolutionary indicators, such as dust temperature, surface density, and MMC mass, and compared their distributions for our different morphological sub-samples (Empty, Isolated, Simple, and Rich). In general, we find that clumps exhibiting an increased level of hierarchical complexity have distributions shifted towards higher dust temperatures, surface densities and MMC masses. The KS-tests performed in Section IV.3 confirm that our morphological sub-samples have distinct distributions in most of the properties surveyed, with significant ($p<0.05$) evidence they do not draw from the same underlying distribution. ALMAGAL targets from the Rich sub-sample show the most structural hierarchy and have an L/M ratio distribution that heavily skews towards higher L/M, with a median value of 20.7 L/M. We consider clumps with L/M >10>10 L/M to be at a later stage of evolution, undergoing active star formation with the possible presence of Hii regions (Cesaroni et al., 2025).
@@
Additionally, as seen in Figure 5, the median MMC masses, dust temperatures and deconvolved surface densities for each clump increases as regions exhibit more morphological complexity. This may be due to the increasing fragmentation of the clump as it evolves (Coletta et al., 2025). Additionally, the amount of material that has been funneled down to the $\sim$1000s au scales that we can observe in ALMAGAL should be higher for more evolved regions as they have had more time to accrete material from the intra-clump gas reservoir, which may contribute to the increased richness of structure that we see towards those targets.
V.2 Core Mass Accretion at Different Evolutionary Stages
@@
We investigate the correlation between the MMC mass in a clump and its surrounding surface density, $\Sigma_{\text{surr}}$ (see Section III.3 for details) in the left panel of Figure 8. We see that the MMC mass is directly related to the surrounding surface density in its environment, with MMCs existing only in proportionally dense environments. Different evolutionary stages follow a similar linear trend, with clumps at later evolutionary stages (higher $L/M$) exhibiting more massive MMCs in higher density environments.
@@
We extend this analysis of comparing the core mass with its surrounding surface density, $\Sigma_{\text{surr}}$, to all sources in the ALMAGAL core catalog for which there is a corresponding dendrogram structure, and present these results in the right panel of Figure 8. When we include the entire core population, we find that the core mass has a looser correlation with its surrounding surface density than when we consider just the MMCs. For all three evolutionary populations shown in Figure 8, the distribution of $M_{\text{core}}$ vs $\Sigma_{\text{surr}}$ covers a larger parameter space than the distribution of $M_{\text{MMC}}$ vs $\Sigma_{\text{surr}}$, with the inclusion of lower mass cores located in higher surrounding surface density environments. This spread in the distribution is exaggerated for more evolved regions, with the extent of the outermost contours in Figure 8 increasing by a factor of $\sim 2$ for both $M_{\text{core}}$ and $\Sigma_{\text{surr}}$ between the $L/M\leq 1~L_{\odot}/M_{\odot}$ and $L/M>10~L_{\odot}/M_{\odot}$ populations.
@@
One interpretation for this trend is that cores competitively accrete mass from an intra-clump reservoir of gas, with some cores accreting mass more efficiently than others depending on their location within the clump (Bonnell et al., 2001; Bonnell and Bate, 2006). In this competitive accretion scenario, cores that are formed later, or in a location outside of the central region of the clump will have a slower mass accretion rate, even as more gas is funneled down from the larger scale reservoir. If we examine the right panel of Figure 8 in the context of core mass accretion throughout the evolution of a clump, we see that cores in clumps at early stages of their evolution have a tighter relationship between $M_{\text{core}}$ and $\Sigma_{\text{surr}}$. As this clump evolves, certain cores will continue to accrete material and maintain a $\Sigma_{\text{surr}}$ that is proportional to their mass. This is especially true for the MMCs, which have masses that are proportional to their surrounding surface densities, regardless of clump evolutionary stage. However in the same amount of time, other member cores will accrete less mass, while a steady stream of material from the intra-clump reservoir continues to increase their surrounding surface densities, which would reproduce the weaker $M_{\text{core}}$ vs. $\Sigma_{\text{surr}}$ correlation that we observe for clumps with higher $L/M$ ratios.
@@
V.3 Sources of Uncertainty
We acknowledge that there are multiple sources of uncertainty in our measurements, in particular our mass estimations as they are highly reliant on our dust temperature assumptions. However, as discussed in Appendix C, if we instead assume the dust temperature of the structure is equal to the dust temperature of its host clump, $T_{\text{clump}}$, we find that the relationship shown in Figure 8 still holds, but exhibits a systematic shift towards higher surrounding surface densities. This makes sense as $T_{\text{clump}}$ is often less than the assumed core-scale temperature from the model we use in Section III.3, which would result in higher masses, and consequently, surface densities for all structures.
@@ V.3 Sources of Uncertainty
It is also possible that our results were impacted by our choice of dendrogram parameters. However, after generating several other versions of the catalog, we did not find any significant changes in our results. We provide more information on the range of parameters investigated and one explanatory figure in Appendix B.
We also explored how missing flux caused by short-spacings could impact our results. In Appendix D, we discuss single-dish combination using continuum observations from the ATLASGAL Survey (Schuller, F. et al., 2009), and quantitatively determine that for most structures, mass calculations differ by less than 10% when using the single-dish combined data instead. Additionally, since feathered data includes more extended emission, our dendrogram structures would have higher measured surface densities while the core masses would remain largely the same. In this scenario, the range of local surface densities calculated from dendrogram structures would increase, causing a larger spread along the $x$-axis of Figure 8, which is still consistent with the results we present in Section V.2. Given this information, and the fact that only 856 out of 900 targets have overlap with the ATLASGAL survey, we ultimately decided against using single-dish combined images in our analysis.
We also consider how the mass estimates of cores from the ALMAGAL core catalog may ultimately affect the results we present in Section V.2. We note that the presence of lower mass cores in more evolved regions may in part be due to the temperature model used (see Section III.3), resulting in the underestimation of mass for more recently formed cores that are at a lower temperature than would be assumed based on the evolutionary stage of the clump. However, this same temperature model would likely also underestimate the cores surrounding surface density, so we consider the presence of lower mass cores at higher $\Sigma_{\text{surr}}$ to be a robust trend despite our temperature assumptions. Although the near sample targets have more low mass core detections than the far sample due to the difference in mass sensitivity, both groups sample a comparable number of clumps in the different evolutionary bins (see Table 2), so we consider this to have a negligible impact on the results shown in Figure 8.
@@ VI Conclusion
    Using astrodendro, we identify 5160 structures across 904 dense clumps targeted in the ALMAGAL survey. These structures have effective radii spanning 80042000800-42000 au and cover multiple orders of magnitude in mass, from 0.056700.05-670 M.
@@
In this paper, we described the process for generating the ALMAGAL hierachical continuum structure catalog and presented the distribution of the fundamental properties of these structures. We found that the level of fragmented sub-structure in clumps is related to their evolution. Additionally, we used the hierarchical continuum structure to infer the mass evolution of cores in clumps at different evolutionary stages, finding evidence of a clump-fed accretion scenario. In future work, this hierarchical structure catalog will be a useful tool for assessing how properties such as the gas temperature and the abundance of molecular species changes across different spatial scales, providing context for the dynamical evolution of these star forming systems.
VII Acknowledgments